# TinyMyo: a Tiny Foundation Model for Flexible EMG Signal Processing at the Edge

### v1

@@ 3.2.3 Experimental Settings
All downstream experiments are performed with the same hyperparameters, as per Table 4.
Table 4: Hyperparameters for model fine-tuning.
Hyperparameter Value
Batch size 32
Peak LR $5\times 10^{-4}$
Minimum LR $1\times 10^{-5}$
LR scheduler Cosine
Optimizer $(\beta_{1},\beta_{2})$ AdamW (0.9, 0.98)(0.9,\ 0.98)
Weight decay 0.01
Max training epochs * 50
Warm-up epochs 5
Drop-path 0.10
Layer-wise LR decay 0.90
Label smoothing 0.10
LR: learning rate;
* Early stopping is adopted, with patience of 7 epochs (validation set).
We evaluate the downstream tasks under two different settings: (i) Full-supervised (FS), where the model is trained from scratch on the task-specific dataset without relying on pre-trained weights; (ii) Full Finetuning (FT), where the entire model is fine-tuned using layer-wise learning rate decay to avoid catastrophic forgetting [kenneweg_forgetting_22].
@@ 3.3 Model Deployment
We deploy the TinyMyo on the GreenWaves Technologies GAP9 [greenwaves_gap_sdk] processor, a multi-core RISC-V MCU designed for ultra-low-power edge AI. The architecture features a hierarchical memory organization consisting of a large off-chip L3 (HyperRAM), a 1.5 MB on-chip L2 (shared memory), and a 128 kB L1 (Tightly Coupled Data Memory) shared among a cluster of 8 compute cores and a dedicated cluster controller.
@@
Unlike traditional MCU inference engines that require the full model to reside in on-chip memory, our approach treats L2 as a streaming buffer. We implement a multi-level tiling strategy:
@@
We apply INT8 quantization to convert all weights and activations from FP32 to 8-bit fixed-point representations. All intermediate tensors, including those in MHSA (with integer softmax) and GELU/LayerNorm, are stored and processed in INT8 format, with only the LayerNorm scale and shift parameters ($\gamma$ and $\beta$) retained in FP32. The only remaining FP32 operation is the final classification layer, which consumes INT8 inputs and weights but accumulates to FP32 logits to enable a direct comparison with the original FP32 baseline; in principle, this layer could also be executed entirely in INT8.
@@
To maintain numerical fidelity while avoiding costly floating-point operations, all transformer primitives are implemented in fixed-point arithmetic[kim_2021_bert]. We use integer-only variants of softmax (i-Softmax), layer normalization (i-LayerNorm), and GELU activation (i-GELU), which leverage lookup tables and multiplicative inverses to approximate their floating-point counterparts with minimal error.
As proof of concept, we deploy the Hand Gesture Classification downstream task on the NinaPro DB5 dataset.
Table 5: FLOPs breakdown
Experiment FLOPs
Model Pretraining (Encoder + Decoder) 4.001 G
Ninapro DB5 4.813 G
EPN-612 1.916 G
UCI-EMG 1.916 G
Generic Neuromotor 13.6315 G
Ninapro DB8 4.852 G
Speech Production 1.0956 G
Speech Recognition 1.0921 G
@@
PhysioWave-Small [chen_2025_physiowave] 5M 84.78 72.54 N.A. 93.12 93.40 N.A. 90.35 89.51 N.A.
@@
Notes: PhysioWave [chen_2025_physiowave] use windows of $\approx 1953\text{ms}$. EMG data are upsampled at $2\text{kHz}$ and a fixed window size of 1024 samples is used. WaveFormer [chen_2025_waveformer] uses a fixed window of 1024 samples; no information is available on whether resampling is applied.
@@
4 Results
4.1 Pre-training
Figure 2: Example of reconstruction for a EMG window by the proposed Foundation model. Grey areas correspond to the masked signal regions, white areas to the unmasked.
Figure 2 shows an example of the reconstruction of the EMG signal performed by the model.
With a parameter count of only 3.6 million (with only 0.1% of parameters for the decoder), in line with previous works [chen_2025_waveformer], our model demonstrates that high reconstruction fidelity can be maintained even under a strict parameter budget and low computational requirements A breakdown of the floating-point operations (FLOPs) is available in Table 5.
@@ 1. Hand Gesture Classification
Table 6 reports the results obtained on the Ninapro DB5[ninapro_db5], EPN-612 [benalcazar_emg-epn-612_2020] and the UC Irvine[emg_data_for_gestures_481] datasets, offering a comparison with previous works based on foundation models.
On the NinaPro DB5 dataset, the fine-tuned model trained with a window size of $200\,\text{ms}$ achieves state-of-the-art performance, with an accuracy of $89.41\pm 0.16\,\%$ and an F1-score of $77.97\pm 0.59\,\%$. Fine-tuning the pre-trained model yields an improvement of approximately $7.5\%$ in accuracy compared to training from scratch. Moreover, using a window size of $200\,\text{ms}$ substantially improves performance over $1000\,\text{ms}$ (around $10\%$ for the fine-tuned model). This effect is consistent with the acquisition protocol, which comprises $52$ different movements. Since, during data pre-processing, we did not remove the transition periods between gestures in order to emulate real-life use cases, using shorter windows reduces the likelihood that segments overlapping these transitions are assigned incorrect gesture labels, thereby mitigating label noise. It must be noted that this dataset exhibits a substantial imbalance toward the rest class. To ensure a fair comparison with prior work, we did not apply any rebalancing techniques during training or evaluation. We assume that earlier studies reported the micro accuracy, as no alternative definition was explicitly stated. When computing the macro-averaged accuracy score, the accuracy decreases to $75.80\pm 0.62\,\%$, which reflects the imbalance in the dataset.
@@
LSTM, Kaifosh [generic_neuromotor_interface_2025] (a) 6.4M 0.1819 Full sequence
LSTM, Kaifosh [generic_neuromotor_interface_2025] (b) 6.4M 0.1596 Windowed
@@
(a): Full sequence results obtained with methods provided by the authors ([generic_neuromotor_interface_2025]) (b): results obtained with the same model as in a, with the windowed-inference method described in the corresponding Methods section. FS: model trained from scratch, FT: fine-tuning pretrained model
For the Generic Neuromotor Interface - discrete gesture task [generic_neuromotor_interface_2025], we achieve a CLER of $0.144\pm 0.004$ on the model trained from scratch, which compares to the 0.1596 (obtained with our windowed inference method, see Sec. 3.2.2) obtained with the LSTM model proposed in the original work, however with $\approx 44\%$ reduction in model size. In this case, the results achieved with the pre-trained Foundation Model are worse ($0.153\pm 0.006$), which we attribute to the intrinsic nature of the problem: to emulate the sequential processing behaviour of the original stacked LSTMs, we imposed a causal mask on the attention mechanism. This constraint restricts our architecture, initially optimized for bidirectional dependency modelling, to a strictly left-to-right context, leading to a degradation in the effectiveness of the pre-trained weights when the future information on which they were originally trained is no longer available. Table 7 reports a summary of the results.
We note that our work is the first to include this dataset among the downstream tasks and to propose an alternative architecture compared to the one employed in the original work.
@@ 5 DoFs, subject-specific model; ∗∗ 3 DoFs, subject-specific model;
∗∗∗ 5 DoFss, across-subject model;
FS: model trained from scratch, FT: fine-tuning pretrained model
We obtain a MAE of $8.77\pm 0.12^{\circ}$, with a window size of $1000\text{ms}$ . This value is higher than the result reported by Zanghieri et al. [zanghieri_2021_db8] (MAE $6.89^{\circ}$), primarily due to differences in both the training and evaluation protocols. In our setting, the model is trained in an across-subject regime using data aggregated from all participants, while prior work trained subject-specific models, i.e., one model per individual. Learning a single model that generalizes across subjects is substantially more challenging because of inter-subject variability in muscle physiology, electrode placement, and EMG signal morphology.
@@ 3. Speech Recognition
Transformer, Gaddy [gaddy2021improved] CNN + Transformer 54 HiFi-GAN 13.92 DeepSpeech 36%
Diff-ETS, Ren [ren_speech_24] as [gaddy2021improved] + Diffusion Probabilistic Model >>54 (a) HiFi-GAN 13.92 Whisper Medium (EN) 32%
SU-E2S, Scheck. [Scheck2023SUE2S] HuBERT-Soft N.A. HiFi-GAN 13.92 Whisper Medium (EN) 40%
DiffMV-ETS [Scheck2025DiffMVETSDM] DiffMV-ETS $\approx$ 66 (b) BigVGAN 112 Distil-Whisper-v3 N.A.
EMGVox-GAN [SUALIHEEN2025101754] as [gaddy2021improved] as [gaddy2021improved] EMGVox-GAN 12 N.A. 36%
SWS, Lee [lee_speech_2025] as [gaddy2021improved] 54 Sovits-SVC N.A. Whisper Medium (EN) 38%
@@
(a) = estimated. The authors employ the same Transduction model as in Gaddy [gaddy2021improved], followed by a Diffusion Model (including residual blocks, attention and convolutional layers)
(b) = estimated as sum of speaker encoder (6M), prior encoder (30M), and diffusion decoder (30M)
FS: model trained from scratch, FT: fine-tuning pretrained model
Table 9 reports the results for the Speech Synthesis Task.
Our model achieves a WER (FT) of $33.54\pm 1.12\%$, in line with previous works (SoA WER is 32%, achieved by Ren et al. [ren_speech_24]). Compared to previous solutions, our Transduction model offers a significant reduction in parameter size, with 62.5% reduction compared to Scheck et al. [Scheck2023SUE2S] and 91.7% reduction compared to Gaddy et al. [gaddy2021improved].
The reported number of parameters accounts only for Transduction model (encoder and model head). The number of parameters for the pre-trained Hi-Fi Gan vocoder is 13.92M (same as in [gaddy2021improved, ren_speech_24]).
@@
Transformer, Gaddy [gaddy_2022_phdthesis] 54 M 28.8%
MONA, Benster [benster_2024_mona] (a) n.a. 22.2%
MONA LISA, [benster_2024_mona] (a) n.a. 12.2%
@@
(a): Multimodal model
FT: fine-tuning pretrained model
Table 10 reports the results for the Speech Recognition Task. Given the superior performance of the fine-tuned model compared to training from scratch in the Speech Production task, and the identical architecture adopted for the vocoder, in this case, we run the experiments only for the full fine-tuning setting. We achieve a WER of $33.95\%\pm 0.97\%$, higher compared to 28.8% reported of Gaddy et al. [gaddy_2022_phdthesis], however, with a 91.7% reduction in parameter cont.
The MONA LISA model [benster_2024_mona] achieves a WER of 12.2%. However, this performance relies on a multimodal training setup, where both EMG and audio data are jointly used. Moreover, MONA LISA is a large-scale language model, which poses practical limitations for real-time deployment and edge applications due to its substantial computational footprint and memory requirements.
@@ 4.3 Model Deployment
We deploy TinyMyo on the GAP9 microcontroller, as described in Section 3.3. The model comprises 3.56 million parameters, with the majority residing in the 8 transformer blocks. Each block contains a multi-head self-attention (MHSA) layer with 3 heads and a feed-forward network (FFN) with an expansion ratio of 4.
Table 11 reports a computational cost breakdown of the deployed model for one Transformer Block. We can observe that the computational cost is dominated by the MHSA QK/AV operations (40% of total MACs) and the MLP projections (40% of total MACs). In particular, the QK and AV terms scale as $\mathcal{O}(N^{2}d)$ with the sequence length $N=800$, making them the asymptotic computational bottleneck of the Transformer block.
With GAP9 operating at 370 MHz, utilizing all 9 cores in the cluster to execute TinyMyo computations in parallel, we achieve an inference in around $12.2\text{seconds}$ at only $0.44\text{J}$ with an average power-envelope of $36.45\text{mW}$.
Table 11: Computational Cost per Transformer Block
Component MACs % of Total
MHSA Q/K/V projections 88M 15%
MHSA QK scores 123M 20%
MHSA AV context 123M 20%
MHSA output projection 29M 5%
MLP FC1 118M 20%
MLP FC2 118M 20%
Notes: QK scores and AV context is (800 $\times$ 800 $\times$ 3 heads). FC1 and FC2 are $192\to 768$ and $768\to 192$ with $4\times$ MLP expansion factor.
@@ 5 Conclusion
In this work, we presented TinyMyo, the first compact EMG FM capable of broad generalization across diverse downstream tasks and real-time edge deployment on an ultra-low-power microcontroller.
@@
TinyMyo achieves SoA performance across multiple downstream tasks. For gesture recognition, we obtain SoA accuracy on the NinaPro DB5 ($89.41\pm 0.16\,\%$), EPN-612 ($96.74\pm 0.09\,\%$), and UCI-EMG ($97.56\pm 0.32\,\%$) datasets. Furthermore, on the Generic Neuromotor Interface – discrete-gesture dataset, TinyMyo achieves a CLER of $0.153\pm 0.006$ with an approximate $44\,\%$ reduction in model size compared to the original work.
@@
TinyMyo is also the first FM for EMG deployed on the edge, achieving an inference in $12.2\text{s}$ at $0.44\text{J}$, with an average power-envelope of $36.45\text{mW}$.
We identify the following limitations, which will be addressed as part of future work:
@@
  • Despite demonstrating, to the best of our knowledge, the first deployment of an FM for EMG processing on an ultra–low-power microcontroller, the inference time remains too high for seamless use in real-world applications. Future work should therefore focus on optimizing the network architecture to minimize the computation needed for inference. This could be achieved by doing an ablation study on reducing the sequence length, by adopting a cheaper attention mechanism [dimofte_2025_cerebro] or exploring more aggressive quantization techniques (e.g.: sub 8-bit [dong_2023_packqvit]. More efficient architectures, such as state-space models (e.g., MAMBA [gu_2024_mamba], FEMBA [tegon_2025_femba]), whose computational cost scales linearly with the context length, would also be worth exploring. Moreover, distributing the computation over multiple GAP9 nodes or deploying it on more powerful platforms such as Siracusa [prasad_2024_siracusa] is expected to yield substantial speed-ups.
Architectural improvements and refinements in the training pipeline also represent an important future direction. We aim to explore alternative pooling mechanisms (e.g., attention-based) to weight per-channel contribution more precisely.
By open-sourcing111after review our pre-trained models and the architectures used for downstream evaluations, we aim to provide a versatile resource that can accelerate future research and serve as a foundation for the broader EMG community. Our intention is that other researchers will extend this backbone with additional datasets, improved training strategies, and architectural variations, aligning with the community-wide effort recently advocated by Eddy et al. [eddy_2025_raising].

### v2

@@ 3.2.3 Experimental Settings
All downstream experiments use the AdamW optimizer ($\beta_{1}=0.9,\beta_{2}=0.98$) with a batch size of 32, weight decay 0.01, and label smoothing 0.1. We employ a cosine learning rate schedule (50 epochs, 5 warm-up) peaking at $5\times 10^{-4}$ (min $1\times 10^{-5}$) with 0.9 layer-wise decay and 0.1 drop-path. Early stopping is adopted with a patience of 7 epochs.
We evaluate the downstream tasks under two different settings: (i) Full-supervised (FS), where the model is trained from scratch on the task-specific dataset without relying on pre-trained weights; (ii) Full Finetuning (FT), where the entire model is fine-tuned using layer-wise learning rate decay to avoid catastrophic forgetting [28].
@@ 3.3 Model Deployment
We deploy the TinyMyo on the GreenWaves Technologies GAP9 [21] processor, a multi-core RISC-V MCU designed for ultra-low-power edge AI. The architecture features a hierarchical memory organization consisting of a large off-chip L3 (HyperRAM), a 1.5 MB on-chip L2 (shared memory), and a 128 kB L1 (Tightly Coupled Data Memory) shared among a cluster of 8 compute cores and a dedicated cluster controller.
@@
Unlike traditional MCU inference engines that require the full model to reside in on-chip memory, we treat L2 as a streaming buffer. We implement a multi-level tiling strategy:
@@
We apply INT8 quantization to convert all weights and activations from FP32 to 8-bit fixed-point representations. All intermediate tensors, including those in MHSA (with integer softmax) and GELU/LayerNorm, are stored and processed in INT8 format, with only the LayerNorm scale and shift parameters ($\gamma$ and $\beta$) retained in FP32. The only remaining FP32 operation is the final classification layer, which consumes INT8 inputs and weights but accumulates to FP32 logits to enable a direct comparison with the original FP32 baseline; in principle, this layer could also be executed in INT8.
@@
To maintain numerical fidelity while avoiding costly floating-point operations, all transformer primitives are implemented in fixed-point arithmetic[29]. We use integer-only variants of softmax (i-Softmax), layer normalization (i-LayerNorm), and GELU activation (i-GELU), which leverage lookup tables and multiplicative inverses to approximate their floating-point counterparts with minimal error.
As proof of concept, we deploy the Hand Gesture Classification downstream task on the EPN-612 dataset.
4 Results
4.1 Pre-training
Figure 2: Example of reconstruction for an EMG window by TinyMyo. Grey areas: masked regions. White areas: unmasked regions.
Figure 2 shows an example of the reconstruction of the EMG signal performed by the model. With a parameter count of only 3.6 million (with only 0.1% of parameters for the decoder), in line with previous works [11], our model demonstrates that high reconstruction fidelity can be maintained even under a strict parameter budget and low computational requirements In terms of computational complexity, the model requires 4.0 GFLOPs for pre-training. Inference costs vary by downstream task complexity: approximately 1.1 GFLOPs for speech tasks, 1.9 GFLOPs for EPN-612/UCI-EMG, 4.8 GFLOPs for Ninapro datasets, and 13.6 GFLOPs for the Generic Neuromotor.
@@ 4.1 Pre-training
PhysioWave-Small [10] 5M 84.78 72.54 N.A. 93.12 93.40 N.A. 90.35 89.51 N.A.
@@
Notes: PhysioWave [10] use windows of $\approx 1953\text{ms}$. EMG data are upsampled at $2\text{kHz}$ and a fixed window size of 1024 samples is used. WaveFormer [11] uses a fixed window of 1024 samples; no information is available on whether resampling is applied.
@@ 1. Hand Gesture Classification
Table 2 reports the results obtained on the Ninapro DB5, EPN-612, and the UC Irvine datasets, comparing TinyMyo to previous FM works.
On the NinaPro DB5 dataset, the fine-tuned model trained with a window size of $200\,\text{ms}$ achieves state-of-the-art performance, with an accuracy of $89.41\pm 0.16\,\%$ and an F1-score of $77.97\pm 0.59\,\%$. Fine-tuning the pre-trained model yields an improvement of approximately $7.5\%$ in accuracy compared to training from scratch. Moreover, using a window size of $200\,\text{ms}$ substantially improves performance over $1000\,\text{ms}$ (around $10\%$ for the fine-tuned model). This effect is consistent with the acquisition protocol, which comprises $52$ different movements. Since, during data pre-processing, we did not remove the transition periods between gestures in order to emulate real-life use cases, using shorter windows reduces the likelihood that segments overlapping these transitions are assigned incorrect gesture labels, thereby mitigating label noise. It must be noted that this dataset exhibits a substantial imbalance toward the rest class. To ensure a fair comparison with prior work, we did not apply any rebalancing techniques during training or evaluation. Computing a macro-averaged accuracy score reduces the accuracy to $75.80\pm 0.62\,\%$, as expected for an imbalanced dataset.
@@
LSTM, Kaifosh [26] (a) 6.4M 0.1819 Full sequence
LSTM, Kaifosh [26] (b) 6.4M 0.1596 Windowed
@@
(a): Full sequence results obtained with methods provided by the authors ([26]) (b): results obtained with the same model as in a, with the windowed-inference method described in the corresponding Methods section. FS: model trained from scratch, FT: fine-tuning pretrained model
For the Generic Neuromotor Interface - discrete gesture task [26], we achieve a CLER of $0.144\pm 0.004$ on the model trained from scratch, which compares to the 0.1596 (obtained with our windowed inference method, see Sec. 3.2.2) obtained with the LSTM model proposed in the original work, however with $\approx 44\%$ reduction in model size. In this case, the results achieved with the pre-trained FM are worse ($0.153\pm 0.006$), which we attribute to the intrinsic nature of the problem: to emulate the sequential processing behaviour of the original stacked LSTMs, we imposed a causal mask on the attention mechanism. This constraint restricts our architecture, initially optimized for bidirectional dependency modelling, to a strictly left-to-right context, leading to a degradation in the effectiveness of the pre-trained weights when the future information on which they were originally trained is no longer available. Table 3 reports a summary of the results.
We note that our work is the first to include this dataset among the downstream tasks and to propose an alternative architecture compared to the one of the original work.
@@ 5 DoFs, subject-specific model; ∗∗ 3 DoFs, subject-specific model;
∗∗∗ 5 DoFs, across-subject model; FS: trained from scratch; FT: fine-tuning
We obtain a MAE of $8.77\pm 0.12^{\circ}$, with a window size of $1000\text{ms}$. This value is higher than the result of. [58] (MAE $6.89^{\circ}$), primarily due to differences in both the training and evaluation protocols. In our setting, the model is trained in an across-subject regime using data aggregated from all participants, while prior work trained subject-specific models, i.e., one model per individual. Learning a single model that generalizes across subjects is substantially more challenging because of inter-subject variability in muscle physiology, electrode placement, and EMG signal morphology.
@@ 3. Speech Recognition
Transformer, Gaddy [17] CNN + Transformer 54 HiFi-GAN 13.92 DeepSpeech 36%
Diff-ETS, Ren [42] as [17] + Diffusion Probabilistic Model >>54 (a) HiFi-GAN 13.92 Whisper Medium (EN) 32%
SU-E2S, Scheck. [45] HuBERT-Soft N.A. HiFi-GAN 13.92 Whisper Medium (EN) 40%
DiffMV-ETS [44] DiffMV-ETS $\approx$ 66 (b) BigVGAN 112 Distil-Whisper-v3 N.A.
EMGVox-GAN [48] as [17] as [17] EMGVox-GAN 12 N.A. 36%
SWS, Lee [33] as [17] 54 Sovits-SVC N.A. Whisper Medium (EN) 38%
@@
(a) = estimated. The authors employ the same Transduction model as in Gaddy [17], followed by a Diffusion Model (including residual blocks, attention and convolutional layers). (b) = estimated as sum of speaker encoder (6M), prior encoder (30M), and diffusion decoder (30M)
FS: model trained from scratch. FT: fine-tuning pretrained model
Table 5 reports the results for the Speech Synthesis Task. We achieve a WER (FT) of $33.54\pm 1.12\%$, in line with previous works (SoA WER is 32%, achieved by Ren et al. [42]). Compared to previous solutions, our Transduction model offers a significant reduction in parameter size, with 62.5% reduction compared to Scheck et al. [45] and 91.7% reduction compared to Gaddy et al. [17].
The reported number of parameters accounts only for Transduction model (encoder and model head). The number of parameters for the pre-trained Hi-Fi Gan vocoder is 13.92M (same as in [17, 42]).
@@
Transformer, Gaddy [18] 54 M 28.8%
MONA, Benster [7] (a) n.a. 22.2%
MONA LISA, [7] (a) n.a. 12.2%
@@
(a): Multimodal model. FT: fine-tuning pretrained model
Table 6 reports the results for the Speech Recognition Task. Given the superior performance of the fine-tuned model compared to training from scratch in the Speech Production task, and the identical architecture adopted for the vocoder, in this case, we run the experiments only for the full fine-tuning setting. We achieve a WER of $33.95\%\pm 0.97\%$, higher compared to 28.8% reported of Gaddy et al. [18], however, with a 91.7% reduction in parameter cont.
The MONA LISA model [7] achieves a WER of 12.2%. However, this performance relies on a multimodal training setup, where both EMG and audio data are jointly used. Moreover, MONA LISA is a large-scale language model, which poses practical limitations for real-time deployment and edge applications due to its substantial computational footprint and memory requirements.
@@ 4.3 Model Deployment
We deploy TinyMyo on the GAP9 microcontroller (MCU), as described in Section 3.3. We achieve an inference time of $0.785~\text{seconds}$ (291.3 M cycles), with energy consumption of $44.90~\text{mJ}$ and average power-envelope of $57.18~\text{mW}$, with GAP9 operating at 370 MHz and using Our custom deployment achieves a high compute utilization of 88.4%; DMA load/store operations account for only $0.4\%$ of cycles, while idle time and overhead are limited to $11.2\%$. Crucially, we achieve a DMA/Compute overlap of $99.4\%$, effectively hiding data transfer latencies.
The model comprises 3.56M parameters, with the majority residing in the 8 transformer blocks. Each block contains a multi-head self-attention (MHSA) layer with 3 heads and a feed-forward network (FFN) with an expansion ratio of 4.
Table 7 summarize the runtime breakdown. The overall utilization is high with limited DMA overhead. The total cycle count is dominated by the eight Transformer blocks. The computational cost for one Transformer block (Table 7) is dominated by the Multi-Head Attention (51.9% of total cycles). In particular, the Q/K/V Projections, QK Attention Scores, and AV terms scale as $\mathcal{O}(N^{2}d)$ with the sequence length $N=400$, making them the asymptotic computational bottleneck of the Transformer block. Crucially, the INT8 kernels reach strong throughput, with several attention and MLP sub-operations exceeding 991010 MACs/cycle (Table 7 The MLP blocks in the model head account for 33.7% of the total cycles. )
Table 7: Sub-operation Breakdown for one transformer block
Operation Cycles (M) % MACs (M) MACs/Cyc
Pre-Attention Norm 2.2 6.2% 0.4 0.17
Multi-Head Attention 18.6 51.9% 120.4 6.46
   Q/K/V Projections 5.7 15.9% 44.2 7.76
   Permute/Reshape 0.1 <<0.1%
   QKT Attention Scores 3.4 9.4% 30.7 9.08
   Softmax Normalization 4.5 12.6%
[…]
   Output Projection 1.9 5.3% 14.7 7.75
Pre-MLP Norm 2.1 5.9% 0.4 0.18
[…]
GELU Activation 0.8 2.3% 3.1 3.73
[…]
Total 35.9 100% 242.2 6.74
@@ 5 Conclusion
We presented TinyMyo, the first compact EMG FM capable of broad generalization across diverse downstream tasks and real-time edge deployment on an ultra-low-power MCU.
@@
TinyMyo achieves SoA performance across multiple downstream tasks. For gesture recognition, we obtain SoA accuracy on the NinaPro DB5 ($89.41\pm 0.16\,\%$), EPN-612 ($96.74\pm 0.09\,\%$), and UCI-EMG ($97.56\pm 0.32\,\%$) datasets. Furthermore, on the Generic Neuromotor Interface – discrete-gesture dataset, TinyMyo achieves a CLER of $0.153\pm 0.006$ with an approximate $44\,\%$ reduction in model size compared to SoA.
@@
TinyMyo is also the first FM for EMG deployed on the edge, running in only $0.785~\text{s}$, with an energy of $44.91~\text{mJ}$ and an average power-envelope of $57.18~\text{mW}$.
Future work will focus on the following limitations:
@@
  • Despite demonstrating the first deployment of an FM for EMG on an ultra–low-power MCU, the inference time is still too high for full real-time operation in low-latency closed loop applications. Future work will focus on optimizing the network architecture to minimize the computation needed for inference. This could be achieved by doing an ablation study on reducing the sequence length, by adopting a cheaper attention mechanism [12], by exploring more aggressive quantization techniques (e.g.: sub 8-bit [14], and by evaluating state-space models [22, 49], whose computational cost scales linearly with the context length. Also, distributing the computation over multiple GAP9 nodes or deploying it on more powerful platforms [40] would yield substantial speed-ups.
By open-sourcing [^1] https://github.com/pulp-bio/BioFoundation our pre-trained models and the architectures used for downstream evaluations, we aim to provide a versatile resource that can accelerate future research and serve as a foundation for the broader EMG community. Our intention is that other researchers will extend this backbone with additional datasets, improved training strategies, and architectural variations.